If the surface area of a ball is 36 pi cm squared, what is the volume of the ball
Answer by TimothyLamb(4379)

a = 4 pi r^2
v = 4/3 pi r^3
---
36 pi = 4 pi r^2
36 = 4 r^2
r^2 = 36/4
---
r = 6/2
r = 3
---
v = 4/3 pi r^3
v = 4/3 pi 3^3
---
v = 113.09734 cu.cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php